# Decentralized Aerial Manipulation of a Cable-Suspended Load using Multi-Agent Reinforcement Learning

### v2

@@ 5 Conclusion
We introduced a decentralized method using MARL that allows for full-pose control of a cable-suspended load using three MAVs without any inter-MAV communication or neighboring MAV information. The policy is computationally tractable and executes entirely onboard. We proposed a novel action space of accelerations and body rates (ACCBR) along with a robust low-level controller and showcase zero-shot transfer from simulation to real-world deployment. Extensive testing with real MAVs shows that the setpoint tracking performance of our method is comparable to that of the state-of-the-art centralized NMPC [6], despite being fully decentralized and having significantly lower computation time. Our method demonstrates robustness against unknown disturbances, heterogeneous agents, and even the complete in-flight failure of one MAV. We attribute this resilience to two key factors: 1) closed-loop reference tracking by the low-level controller, which maintains stability despite perturbations, 2) decentralized policy independence, where local agents operate without dependence on neighboring states, preventing cascading failures. Our work shows promising results to enable scalable and robust cooperative aerial manipulation with minimal onboard sensing and no internal communications required.
@@ A.1 Experimental setup
Real-world evaluation setup We evaluate our method in real-world experiments. Our experiment includes 3 MAVs built based on the Agilicious [48] flight stack. Each MAV is connected to a basket-shaped payload with 1-meter cables at three distinct locations. The MAVs weigh 0.6kg, and the payload weighs 1.4 kg. We conduct the experiment in an indoor flight space with motion capture systems. We attach motion capture markers to the MAVs and the payload to measure their positions and orientations and distribute them to each MAV through ROS at 100 Hz. The trained policy and low-level controllers are deployed onboard each MAV. The policy is inferred at 100 Hz to send acceleration and body-rate commands. The low-level controller is executed at 300 Hz to generate rotor speed commands.
@@ A.4 Trajectory tracking
Although our method is not trained for trajectory tracking, we evaluate its trajectory tracking capabilities against that of the centralized NMPC [6] in Figure 11. The reference trajectory is a figure-eight trajectory with a maximum velocity of 1 m/s and a maximum acceleration of 0.5 m/s2. It is worth noting that our method only considers the reference pose information, while the NMPC also takes velocity information from the reference trajectory into account. For future specialized trajectory tasks, incorporating higher-order information such as velocity, as well as future reference points [49] into the observations would significantly improve tracking performance and make for a fairer comparison. Nonetheless, our method is able to successfully track the figure-eight trajectory, albeit with a high tracking error. Our method achieves positional and attitude RMSEs of 0.82 m (vs 0.10 m), and 18.22 degrees (vs 4.80 degrees).
Figure 11: Comparison of our method, which is not trained for trajectory tracking, against the centralized NMPC in [6]. Left: top view of the flight path of the center of mass of the load while tracking a figure-eight trajectory with a maximum velocity of 1 m/s and maximum acceleration of 0.5 m/s2. Right: position (top) and attitude (bottom) tracking errors time series.
@@ A.5 Performance without centralized critic
To assess the impact of using a centralized critic with access to privileged global state information, we compare its performance against a policy trained with a shared local critic. The local critic has access only to local observations, which are the same as those available to the actor. The training curves in Figure 12 show that the setup with the local critic fails to converge to the same performance as with the centralized critic, and even collapses at the end. Specifically, the policy with the local critic fails to learn the position and orientation rewards effectively. We hypothesize that access to global state information allows the centralized critic to produce more accurate value estimates, which can indirectly support more effective credit assignment during learning [50], thereby improving task performance.
@@ A.8 Training configuration
The inputs to the network are normalized stacked observation histories with history size $H=3$. We also implement a form of advantage filtering [51] where $50\%$ of the samples with the lowest advantage magnitude are dropped. This approach prioritizes learning from the most informative state transitions—specifically the underexplored extremes of the data distribution where actions have a clearly better or worse outcome—thereby improving data efficiency during training. For a complete overview of the network and agent parameters, we refer the readers to Table 5.
For setups with more than 3 MAVs, the mass of the load is sampled from a uniform distribution between 1.0 and 1.8 kg (the mass of the real payload is 1.4 kg). For the 3-MAV setup, the cables are modeled as rigid rods of 1 meter in length, connected to both the payload and the MAVs via ball joints. When using more than 3 MAVs, the system becomes overconstrained, which can lead to cable slack [12]. To address this, the cables are instead modeled as three rigid segments linked by ball joints.
@@
Hyperparameters of MAPPO The hyperparameters of MAPPO are shown in table 5. The names of the parameters are based on the SKRL [44] learning library.
@@
Table 5: MAPPO hyperparameters based on SKRL [44] learning library

### v3

@@ 5 Conclusion
We introduced a decentralized method using MARL that allows for full-pose control of a cable-suspended load using three MAVs without any inter-MAV communication or neighboring MAV information. The policy is computationally tractable and executes entirely onboard. We proposed a novel action space of accelerations and body rates (ACCBR) along with a robust low-level controller and showcase zero-shot transfer from simulation to real-world deployment. Extensive testing with real MAVs shows that the setpoint tracking performance of our method is comparable to that of the state-of-the-art centralized NMPC [sun2025agilecooperativeaerialmanipulation], despite being fully decentralized and having significantly lower computation time. Our method demonstrates robustness against unknown disturbances, heterogeneous agents, and even the complete in-flight failure of one MAV. We attribute this resilience to two key factors: 1) closed-loop reference tracking by the low-level controller, which maintains stability despite perturbations, 2) decentralized policy independence, where local agents operate without dependence on neighboring states, preventing cascading failures. Our work shows promising results to enable scalable and robust cooperative aerial manipulation with minimal onboard sensing and no internal communications required.
@@ A.1 Experimental setup
Real-world evaluation setup We evaluate our method in real-world experiments. Our experiment includes 3 MAVs built based on the Agilicious [foehn_agilicious_2022] flight stack. Each MAV is connected to a basket-shaped payload with 1-meter cables at three distinct locations. The MAVs weigh 0.6kg, and the payload weighs 1.4 kg. We conduct the experiment in an indoor flight space with motion capture systems. We attach motion capture markers to the MAVs and the payload to measure their positions and orientations and distribute them to each MAV through ROS at 100 Hz. The trained policy and low-level controllers are deployed onboard each MAV. The policy is inferred at 100 Hz to send acceleration and body-rate commands. The low-level controller is executed at 300 Hz to generate rotor speed commands.
@@ A.4 Trajectory tracking
Although our method is not trained for trajectory tracking, we evaluate its trajectory tracking capabilities against that of the centralized NMPC [sun2025agilecooperativeaerialmanipulation] in Figure 11. The reference trajectory is a figure-eight trajectory with a maximum velocity of 1 m/s and a maximum acceleration of 0.5 m/s2. It is worth noting that our method only considers the reference pose information, while the NMPC also takes velocity information from the reference trajectory into account. For future specialized trajectory tasks, incorporating higher-order information such as velocity, as well as future reference points [mayer2021proximal] into the observations would significantly improve tracking performance and make for a fairer comparison. Nonetheless, our method is able to successfully track the figure-eight trajectory, albeit with a high tracking error. Our method achieves positional and attitude RMSEs of 0.82 m (vs 0.10 m), and 18.22 degrees (vs 4.80 degrees).
Figure 11: Comparison of our method, which is not trained for trajectory tracking, against the centralized NMPC in [sun2025agilecooperativeaerialmanipulation]. Left: top view of the flight path of the center of mass of the load while tracking a figure-eight trajectory with a maximum velocity of 1 m/s and maximum acceleration of 0.5 m/s2. Right: position (top) and attitude (bottom) tracking errors time series.
@@ A.5 Performance without centralized critic
To assess the impact of using a centralized critic with access to privileged global state information, we compare its performance against a policy trained with a shared local critic. The local critic has access only to local observations, which are the same as those available to the actor. The training curves in Figure 12 show that the setup with the local critic fails to converge to the same performance as with the centralized critic, and even collapses at the end. Specifically, the policy with the local critic fails to learn the position and orientation rewards effectively. We hypothesize that access to global state information allows the centralized critic to produce more accurate value estimates, which can indirectly support more effective credit assignment during learning [foerster2018counterfactual], thereby improving task performance.
@@ A.8 Training configuration
The inputs to the network are normalized stacked observation histories with history size $H=3$. We also implement a form of advantage filtering [cusumano2025robust] where $50\%$ of the samples with the lowest advantage magnitude are dropped. This approach prioritizes learning from the most informative state transitions—specifically the underexplored extremes of the data distribution where actions have a clearly better or worse outcome—thereby improving data efficiency during training. For a complete overview of the network and agent parameters, we refer the readers to Table 5.
For setups with more than 3 MAVs, the mass of the load is sampled from a uniform distribution between 1.0 and 1.8 kg (the mass of the real payload is 1.4 kg). For the 3-MAV setup, the cables are modeled as rigid rods of 1 meter in length, connected to both the payload and the MAVs via ball joints. When using more than 3 MAVs, the system becomes overconstrained, which can lead to cable slack [fink_planning_2011]. To address this, the cables are instead modeled as three rigid segments linked by ball joints.
@@
Hyperparameters of MAPPO The hyperparameters of MAPPO are shown in table 5. The names of the parameters are based on the SKRL [serrano2023skrl] learning library.
@@
Table 5: MAPPO hyperparameters based on SKRL [serrano2023skrl] learning library